This morning you purchased a stock that just paid an annual dividend of $3.10 per share. You require a return of 9.2 percent and the dividend will increase at an annual growth rate of 4 percent. If you sell this stock in three years, what will your capital gain be?

$4.00

$5.87

$10.13

$7.74

$2.48
Answer #1

SOLUTION :


Current dividend, D0 = 3.10 ($)

Dividend growth rate, g = 4% = 0.04

Market rate of return, r = 9.2 % = 0.092


So, as per Dividend constant growth Discount Model :


Expected price currently, P0 

= D0(1+g) / (r - g) 

= 3.10(1 + 0.04) / (0.092 - 0.04)

= 62 ($).


Price expected after 3 years, P3 = 62(1 + 0.04)^3 = 69.74 ($)


So, capital gain, if stock is sold in 3 years 

= 69.74 - 62

= 7.74 ($) (ANSWER).



answered by: Tulsiram Garg